# Atherosclerosis through Hierarchical Explainable Neural Network Analysis

### v1

@@ Abstract
In this work, we study the problem pertaining to personalized classification of subclinical atherosclerosis by developing a hierarchical graph neural network framework to leverage two characteristic modalities of a patient: clinical features within the context of the cohort, and molecular data unique to individual patients. Current graph-based methods for disease classification detect patient-specific molecular fingerprints, but lack consistency and comprehension regarding cohort-wide features, which are an essential requirement for understanding pathogenic phenotypes across diverse atherosclerotic trajectories. Furthermore, understanding patient subtypes often considers clinical feature similarity in isolation, without integration of shared pathogenic interdependencies among patients. To address these challenges, we introduce ATHENA: Atherosclerosis Through Hierarchical Explainable Neural Network Analysis, which constructs a novel hierarchical network representation through integrated modality learning; subsequently, it optimizes learned patient-specific molecular fingerprints that reflect individual omics data, enforcing consistency with cohort-wide patterns. With a primary clinical dataset of 391 patients [27], their respective transcriptomics signatures, as well as their STRING PPI profiles, we demonstrate that this heterogeneous alignment of clinical features with molecular interaction patterns has significantly boosted subclinical atherosclerosis classification performance across various baselines by up to 13% in area under the receiver operating curve (AUC) and 20% in F1 score. We further validated ATHENA on a secondary clinical dataset on atherosclerosis [30]. Taken together, ATHENA enables mechanistically-informed patient subtype discovery through explainable AI (XAI)-driven subnetwork clustering; this novel integration framework strengthens personalized intervention strategies, thereby improving the prediction of atherosclerotic disease progression and management of their clinical actionable outcomes.
@@
Keywords: Multi-modal Biomedical Data, Atherosclerosis, Graph Deep Learning, Transcriptomics, Explainable AI, Hierarchical Networks, Patient Subtype, Patient Classification, Machine Learning
1Medical Informatics HA,
2Department of Physiology,
3Department of Medicine / Cardiology Division,
4Department of Computer Science,
University of California, Los Angeles
Corresponding author: {irsyadadam, pping38}@g.ucla.edu
1 Introduction
@@
Subclinical atherosclerosis is a widespread yet insidious precursor to life-threatening cardiovascular events [37, 18, 35, 4]. The accumulation of plaque in the arteries and resulting carotid intima-media thickness have been applied as indicators of atherosclerotic pathogenesis [37, 18]. These markers remain imperceptible to patients until occlusion or rupture occurs. Noncontrast CT and other imaging platforms have been widely implemented for the diagnosis and risk stratification of subclinical atherosclerosis [26, 19, 22]. However, these approaches remain operator-dependent and thus, inherently inconsistent, limiting their applications to clinical care [21]. Given the silent nature of this disease, early detection of plasma molecular markers is imperative to guiding tailored intervention before irreversible damage ensues. Molecular profiling is valuable in quantifying patient-specific atherosclerotic burden, helping stage or subtype a patient. However, leveraging specific omics data (e.g. transcriptomics) to better understand clinical subtypes of atherosclerosis remains a challenge in the field.
To leverage the benefits of molecular signatures in phenotypic diagnosis, investigations in omics-based classification techniques have emerged for stratifying disease risk and elucidating underlying drivers of pathogenesis. Traditional approaches often begin with differential expression analysis or penalized regression to identify a subset of features that are correlative of a diagnosis [14]. Supervised classifiers such as support vector machines (SVMs), random forests, and elastic-net logistic regression have then been trained on the feature subset to differentiate and classify phenotypes [8, 17]. Due to the inherent noisy and nonlinear nature of expression data, recent deep-learning methods employ autoencoders or multi-layer perceptrons to learn latent representation of molecular profiles to generate a more robust and effective differentiation [49]. However, many models treat each molecular profile as an unstructured vector and ignore the underlying mechanistic insights reflected in molecular interactions of cellular networks and pathway topologies. We address this challenge by applying graph deep learning models that have been shown to effectively integrate and identify these molecular interactions towards a specific phenotype. Briefly, by embedding molecules as nodes within biological networks and using graph convolutional networks (GCNs) and graph attention networks (GATs) to propagate molecular expression signals between edges, we achieved a representation reflecting both the omics profile and its underlying pathogenesis of the clinical sub-phenotypes [42, 34, 41, 23, 28].
Recent advancements in graph-supported deep learning has expanded our understanding of cellular pathways and networks with respect to genotypes and function. However, effective translation of this field faces major limitations: first, these workflows are limited to a single homogeneous graph and its subgraph relevant to disease, but cannot model individual patient variabilities and understand the heterogeneity of a cohort. Furthermore, these approaches focus solely on molecular information without considering cohort-wide population phenotypes such as shared comorbidities and demographic similarities, which are crucial for developing risk stratification strategies [50]. Finally, there are challenges in model design to characterize subclasses of patient pathogenesis based on their molecular data, a key component for our understanding of major diseases, e.g., subclinical atherosclerosis [25].
To address these specific limitations, we propose a novel graph-based approach that models two essential aspects of a patient: clinical features and molecular data. We first create a patient similarity network based on their clinical features, thereby modelling cohort-level patterns and discovering clinically relevant subpopulations. By representing each patient as a node and weighting each edge according to their clinical feature-based similarities, the network consists of patient subgroups exhibiting similar conditions [20, 32]. Molecular interaction networks encode mechanistic insights by treating proteins as nodes and their interactions as edges, enabling the identification of disease-relevant subnetworks [43]. We demonstrate that jointly embedding patient-specific molecular networks and cohort-level similarity networks into a unified hierarchical graph has shown a holistic representation of disease phenotypes, which improves the accuracy of subclinical atherosclerosis classification and facilitates robust patient subtyping. By capturing both individual perturbations and cohort-level population patterns, our framework delivers interpretable mechanistic insights to better understand disease pathogenesis and to aid personalized intervention strategies.
In a nutshell, we introduce a hierarchical integration framework and graph-representation learning pipeline that jointly leverages patient-specific molecular networks and cohort-level similarity graphs. We evaluate our approach on a cardiovascular case study utilizing a cohort from Sanchez-Cabo et al. [27] and from Steenman et al. [30], as well as demonstrate that a hierarchical representation for patient stratification outperforms previous graph deep learning approaches for classification. Importantly, we demonstrate the capacity of our approach pertaining to disease subtype identification by clustering the learned protein interactions of patient-specific molecular sub-networks. This approach is novel as it combines the endogenous molecular features unique to each patient with that of clinical features reflecting cohort-level patterns, resulting in early risk stratification and interpretable subtype identification to assist targeted intervention strategies.
@@
Figure 1: Transcriptomics-injected patient-specific PPI networks are embedded within their respective clinical features nodes in a patient similarity network.
2 Problem Statement
  1. 1.
  2. 2.
  3. 3.
3 Methods
@@
Sanchez-Cabo et al. [27] Steenman et al. [30]
@@
Table 1: Characteristics of Network Topologies Supporting Graph Construction Strategies. Illustration of node count, edge count, and average degree for the three graph types, including a graph representing clinical features of patients in a cohort setting (patient similarity network), a graph representing a patient-specific STRING Protein-Protein Interaction (PPI) Network cross-referenced with individualized transcriptomics features, and a hierarchical network (i.e., ATHENA) that integrates the cohort-context patient similarity graph with each patients’ respective PPI network. These network characteristics provide context for graph complexity and connectivity in downstream graph-based analyses.
3.1 Datasets and Preprocessing
We employed a transcriptomics dataset from the Progression of Early Subclinical Atherosclerosis (PESA) study [27]. Sanchez-Cabo et al. reported whole blood transcriptomics data from 391 patients. In this study, subclinical atherosclerosis burden was assessed using vascular ultrasound and coronary artery calcification (CAC) scores, enabling the identification of subclinical phenotypic associations, in which molecular profiles across different stages of disease development were measured using RNA sequencing (RNA-seq) of polyadenylated RNA extracted from peripheral blood samples. From these CAC scores, a clinician assigned each patient one of four PESA scores from the ultrasound study: None, Generalized, Intermediate, and Focal. For our analysis, we utilized the plasma transcriptomics assay data, comprising measurements for 12,061 genes per patient, and patient clinical features. To validate our analysis, we also employed another transcriptomics dataset with 104 patients from Steenman et al. [30], which includes RNA sequencing data from human arterial tissue samples collected post-mortem.
For preprocessing transcriptomics, we utilized a variance threshold of 0.1, followed by an iterative one-way analysis-of-variance (ANOVA) with a Bonferroni-adjusted p-value threshold of 0.01 to filter out transcriptomics features with minimal variability across patients. The selected transcriptomics features establish a baseline for interpreting differential gene expression across the four PESA scores. Next, we cross-referenced the genes from the PESA study to the STRING PPI database [31] and applied a high-confidence threshold of 0.7 to remove statistically insignificant associations. Our final transcriptomics dataset consisted of 645 protein-coding genes with 5,327 corresponding interactions from STRING. Using the same preprocessing steps for our second dataset, we obtained a final transcriptomics dataset comprising 327 protein-coding genes and 3,229 corresponding interactions from STRING.
3.2 Hierarchical Graph Construction
@@
where $\sigma$ is a bandwidth parameter controlling the width of the kernel. A $k$-nearest neighbors (kNN) strategy is applied to retain only the top-$k$ most similar patients for each node, resulting in a sparse similarity matrix $S$. For the dataset from Sanchez et al. [27], we selected $k=50$, and for Steenman et al. [30], $k=20$.
@@
Figure 2: Pipeline for multi-modal hierarchical graph deep learning for patient classification. Graph encoders are applied to each patient’s PPI network to generate feature representations that capture both structural and omics-informed embeddings. These embeddings serve as inputs to every patient node in the patient similarity network, which is then processed using a GNN to classify atherosclerotic burden. An explainable AI (XAI) module is linked to the GNN to extract subgraphs from the patient-specific PPI networks, highlighting potential patient-level biomarkers contributing to the predictions.
3.3 Model Architecture and Subtype Characterization
@@
$\mathbf{h}_{v}^{(\ell)}=\sigma\left(\mathbf{W}^{(\ell)}\cdot\text{AGGREGATE}^{(\ell)}\left(\left\{\mathbf{h}_{u}^{(\ell-1)}:u\in\mathcal{N}(v)\cup\{v\}\right\}\right)\right),$ (1)
@@
$\mathbf{z}_{i}=\phi_{\text{PPI}}(\mathcal{G}_{i})=\text{READOUT}\left(\left\{\mathbf{h}_{v}^{(L)}:v\in\mathcal{V}_{i}\right\}\right).$ (2)
@@
$\mathbf{h}_{i}^{(\ell)}=\sigma\left(\mathbf{W}^{(\ell)}\cdot\text{AGGREGATE}^{(\ell)}\left(\left\{\mathbf{h}_{j}^{(\ell-1)}:j\in\mathcal{N}(i)\cup\{i\}\right\}\right)\right),$ (3)
@@
$\hat{y}_{i}=\phi_{\text{cohort}}(\mathbf{h}_{i}^{(L^{\prime})}),$ (4)
@@
Finally, an explainable AI (XAI) module extracts subgraphs from the ATHENA’s predictions, revealing patient-specific biomarkers most influential for a patient’s atherosclerotic burden. Specifically, we employ GNNExplainer [46] to extract a subgraph $\mathcal{G}_{i}^{\text{XAI}}\subseteq\mathcal{G}_{i}$ that maximizes the mutual information between the explanation and the prediction $\hat{y}_{i}$. Formally, GNNExplainer optimizes the following objective:
$\max_{\mathcal{G}_{i}^{\text{XAI}},\,X_{i}^{\text{XAI}}}\;\mathbb{MI}\left(\hat{y}_{i};\mathcal{G}_{i}^{\text{XAI}},X_{i}^{\text{XAI}}\right),$ (5)
@@
4 Results
4.1 Evaluation Metrics
For model evaluations, we employed area under the receiver operator curve (AUC) and F1 score [24]. For benchmarking and in our ablation studies, both metrics were calculated across all phenotypes. For evaluating our Hierarchical Network Classification model on specific atherosclerosis phenotypes, only AUC was calculated. In all evaluations, both metrics were calculated across 5-fold stratified cross validation.
4.2 Classification and Profiling Performance

### v2

# ATHENA: Atherosclerosis through Hierarchical Explainable Neural Network Analysis

@@ Abstract.
In this work, we study the problem pertaining to personalized classification of subclinical atherosclerosis by developing a hierarchical graph neural network framework to leverage two characteristic modalities of a patient: clinical features within the context of the cohort, and molecular data unique to individual patients. Current graph-based methods for disease classification detect patient-specific molecular fingerprints, but lack consistency and comprehension regarding cohort-wide features, which are an essential requirement for understanding pathogenic phenotypes across diverse atherosclerotic trajectories. Furthermore, understanding patient subtypes often considers clinical feature similarity in isolation, without integration of shared pathogenic interdependencies among patients. To address these challenges, we introduce ATHENA: Atherosclerosis Through Hierarchical Explainable Neural Network Analysis, which constructs a novel hierarchical network representation through integrated modality learning; subsequently, it optimizes learned patient-specific molecular fingerprints that reflect individual omics data, enforcing consistency with cohort-wide patterns. With a primary clinical dataset of 391 patients (PESA study), their respective transcriptomics signatures, as well as their STRING PPI profiles, we demonstrate that this heterogeneous alignment of clinical features with molecular interaction patterns has significantly boosted subclinical atherosclerosis classification performance across various baselines by up to 13% in area under the receiver operating curve (AUC) and 20% in F1 score. We further validated ATHENA on a secondary clinical dataset on atherosclerosis (by Steenman and Espitia et al.). Taken together, ATHENA enables mechanistically-informed patient subtype discovery through explainable AI (XAI)-driven subnetwork clustering; this novel integration framework strengthens personalized intervention strategies, thereby improving the prediction of atherosclerotic disease progression and management of their clinical actionable outcomes.
@@
Multi-modal Biomedical Data, Atherosclerosis, Graph Deep Learning, Transcriptomics, Explainable AI, Hierarchical Networks, Patient Subtyping, Patient Classification, Machine Learning
copyright: ccjournalyear: 2025doi: 10.1145/3765612.3767226conference: 16th ACM International Conference on Bioinformatics, Computational Biology and Health Informatics; October 11–15, 2025; Philadelphia, PAisbn: 978-1-4503-XXXX-X/2025/11ccs: Applied computing Biological networksccs: Computing methodologies Artificial intelligenceccs: Graph-based representations
1. Introduction
@@
Subclinical atherosclerosis is a widespread yet insidious precursor to life-threatening cardiovascular events (Ullah et al., 2025; Martin et al., 2025; Torrijo-Belanche et al., 2024; Barkas et al., 2024). The accumulation of plaque in the arteries and resulting carotid intima-media thickness have been applied as indicators of atherosclerotic pathogenesis (Ullah et al., 2025; Martin et al., 2025). These markers remain imperceptible to patients until occlusion or rupture occurs. Noncontrast CT and other imaging platforms have been widely implemented for the diagnosis and risk stratification of subclinical atherosclerosis (Ross, 1999; O’leary et al., 1999; Peters et al., 2012). However, these approaches remain operator-dependent and thus, inherently inconsistent, limiting their applications to clinical care (Papageorgiou et al., 2017). Given the silent nature of this disease, early detection of plasma molecular markers is imperative to guiding tailored intervention before irreversible damage ensues. Molecular profiling is valuable in quantifying patient-specific atherosclerotic burden, helping stage or subtype a patient. However, leveraging specific omics data (e.g. transcriptomics) to better understand clinical subtypes of atherosclerosis remains a challenge in the field.
@@
Figure 1. Transcriptomics-injected patient-specific PPI networks are embedded within their respective clinical features nodes in a patient similarity network.
To leverage the benefits of molecular signatures in phenotypic diagnosis, investigations in omics-based classification techniques have emerged for stratifying disease risk and elucidating underlying drivers of pathogenesis. Traditional approaches often begin with differential expression analysis or penalized regression to identify a subset of features that are correlative of a diagnosis (Libby, 2012). Supervised classifiers such as support vector machines (SVMs), random forests, and elastic-net logistic regression have then been trained on the feature subset to differentiate and classify phenotypes (Hansson, 2005; Maiwald et al., 2013). Due to the inherent noisy and nonlinear nature of expression data, recent deep-learning methods employ autoencoders or multi-layer perceptrons to learn latent representation of molecular profiles to generate a more robust and effective differentiation (Zhao et al., 2024). However, many models treat each molecular profile as an unstructured vector and ignore the underlying mechanistic insights reflected in molecular interactions of cellular networks and pathway topologies. We address this challenge by applying graph deep learning models that have been shown to effectively integrate and identify these molecular interactions towards a specific phenotype. Briefly, by embedding molecules as nodes within biological networks and using graph convolutional networks (GCNs) and graph attention networks (GATs) to propagate molecular expression signals between edges, we achieved a representation reflecting both the omics profile and its underlying pathogenesis of the clinical sub-phenotypes (Way and Greene, 2018; Taroni et al., 2018; Wang et al., 2021; Pfeifer et al., 2022; Schulte-Sasse et al., 2021).
Recent advancements in graph-supported deep learning has expanded our understanding of cellular pathways and networks with respect to genotypes and function. However, effective translation of this field faces major limitations: first, these workflows are limited to a single homogeneous graph and its subgraph relevant to disease, but cannot model individual patient variabilities and understand the heterogeneity of a cohort. Furthermore, these approaches focus solely on molecular information without considering cohort-wide population phenotypes such as shared comorbidities and demographic similarities, which are crucial for developing risk stratification strategies (Zhou et al., 2022). Finally, there are challenges in model design to characterize subclasses of patient pathogenesis based on their molecular data, a key component for our understanding of major diseases, e.g., subclinical atherosclerosis (Rajadhyaksha and Chitkara, 2023).
To address these specific limitations, we propose a novel graph-based approach that models two essential aspects of a patient: clinical features and molecular data. We first create a patient similarity network based on their clinical features, thereby modelling cohort-level patterns and discovering clinically relevant subpopulations. By representing each patient as a node and weighting each edge according to their clinical feature-based similarities, the network consists of patient subgroups exhibiting similar conditions (Pai et al., 2019; Tang et al., 2023). Molecular interaction networks encode mechanistic insights by treating proteins as nodes and their interactions as edges, enabling the identification of disease-relevant subnetworks (Xie et al., 2025). We demonstrate that jointly embedding patient-specific molecular networks and cohort-level similarity networks into a unified hierarchical graph has shown a holistic representation of disease phenotypes, which improves the accuracy of subclinical atherosclerosis classification and facilitates robust patient subtyping. By capturing both individual perturbations and cohort-level population patterns, our framework delivers interpretable mechanistic insights to better understand disease pathogenesis and to aid personalized intervention strategies.
In a nutshell, we introduce a hierarchical integration framework and graph‐representation learning pipeline that jointly leverages patient‐specific molecular networks and cohort‐level similarity graphs. We evaluate our approach on a cardiovascular case study utilizing a cohort from Sanchez-Cabo et al. (Sánchez-Cabo et al., 2023) and from Steenman et al. (Steenman et al., 2018), as well as demonstrate that a hierarchical representation for patient stratification outperforms previous graph deep learning approaches for classification. Importantly, we demonstrate the capacity of our approach pertaining to disease subtype identification by clustering the learned protein interactions of patient-specific molecular sub-networks. This approach is novel as it combines the endogenous molecular features unique to each patient with that of clinical features reflecting cohort-level patterns, resulting in early risk stratification and interpretable subtype identification to assist targeted intervention strategies.
@@
Sanchez-Cabo et al. (Sánchez-Cabo et al., 2023) Steenman et al. (Steenman et al., 2018)
@@
Table 1. Characteristics of Network Topologies Supporting Graph Construction Strategies. Illustration of node count, edge count, and average degree for the three graph types, including a graph representing clinical features of patients in a cohort setting (patient similarity network), a graph representing a patient-specific STRING Protein-Protein Interaction (PPI) Network cross-referenced with individualized transcriptomics features, and a hierarchical network (i.e., ATHENA) that integrates the cohort-context patient similarity graph with each patients’ respective PPI network. These network characteristics provide context for graph complexity and connectivity in downstream graph-based analyses.
2. Problem Statement
  1. (1)
  2. (2)
  3. (3)
3. Methods
3.1. Datasets and Preprocessing
We employed a transcriptomics dataset from the Progression of Early Subclinical Atherosclerosis (PESA) study (Sánchez-Cabo et al., 2023). Sanchez-Cabo et al. reported whole blood transcriptomics data from 391 patients. In this study, subclinical atherosclerosis burden was assessed using vascular ultrasound and coronary artery calcification (CAC) scores, enabling the identification of subclinical phenotypic associations, in which molecular profiles across different stages of disease development were measured using RNA sequencing (RNA-seq) of polyadenylated RNA extracted from peripheral blood samples. From these CAC scores, a clinician assigned each patient one of four PESA scores from the ultrasound study: None, Generalized, Intermediate, and Focal. For our analysis, we utilized the plasma transcriptomics assay data, comprising measurements for 12,061 genes per patient, and patient clinical features. To validate our analysis, we also employed another transcriptomics dataset with 104 patients from Steenman et al. (Steenman et al., 2018), which includes RNA sequencing data from human arterial tissue samples collected post-mortem.
For preprocessing transcriptomics, we utilized a variance threshold of 0.1, followed by an iterative one-way analysis-of-variance (ANOVA) with a Bonferroni-adjusted p-value threshold of 0.01 to filter out transcriptomics features with minimal variability across patients. The selected transcriptomics features establish a baseline for interpreting differential gene expression across the four PESA scores. Next, we cross-referenced the genes from the PESA study to the STRING PPI database (Szklarczyk et al., 2025) and applied a high-confidence threshold of 0.7 to remove statistically insignificant associations. Our final transcriptomics dataset consisted of 645 protein-coding genes with 5,327 corresponding interactions from STRING. Using the same preprocessing steps for our second dataset, we obtained a final transcriptomics dataset comprising 327 protein-coding genes and 3,229 corresponding interactions from STRING.
3.2. Hierarchical Graph Construction
@@
Figure 2. Pipeline for multi-modal hierarchical graph deep learning for patient classification. Graph encoders are applied to each patient’s PPI network to generate feature representations that capture both structural and omics-informed embeddings. These embeddings serve as inputs to every patient node in the patient similarity network, which is then processed using a GNN to classify atherosclerotic burden. An explainable AI (XAI) module is linked to the GNN to extract subgraphs from the patient-specific PPI networks, highlighting potential patient-level biomarkers contributing to the predictions.
@@
where $\sigma$ is a bandwidth parameter controlling the width of the kernel. A $k$-nearest neighbors (kNN) strategy is applied to retain only the top-$k$ most similar patients for each node, resulting in a sparse similarity matrix $S$. For the dataset from Sanchez et al. (Sánchez-Cabo et al., 2023), we selected $k=50$, and for Steenman et al. (Steenman et al., 2018), $k=20$.
@@
3.3. Model Architecture and Subtype Characterization
@@
(1) $\mathbf{h}_{v}^{(\ell)}=\sigma\left(\mathbf{W}^{(\ell)}\cdot\text{AGGREGATE}^{(\ell)}\left(\left\{\mathbf{h}_{u}^{(\ell-1)}:u\in\mathcal{N}(v)\cup\{v\}\right\}\right)\right),$
@@
(2) $\mathbf{z}_{i}=\phi_{\text{PPI}}(\mathcal{G}_{i})=\text{READOUT}\left(\left\{\mathbf{h}_{v}^{(L)}:v\in\mathcal{V}_{i}\right\}\right).$
@@
(3) $\mathbf{h}_{i}^{(\ell)}=\sigma\left(\mathbf{W}^{(\ell)}\cdot\text{AGGREGATE}^{(\ell)}\left(\left\{\mathbf{h}_{j}^{(\ell-1)}:j\in\mathcal{N}(i)\cup\{i\}\right\}\right)\right),$
@@
(4) $\hat{y}_{i}=\phi_{\text{cohort}}(\mathbf{h}_{i}^{(L^{\prime})}),$
@@
Finally, an explainable AI (XAI) module extracts subgraphs from the ATHENA’s predictions, revealing patient-specific biomarkers most influential for a patient’s atherosclerotic burden. Specifically, we employ GNNExplainer (Ying et al., 2019) to extract a subgraph $\mathcal{G}_{i}^{\text{XAI}}\subseteq\mathcal{G}_{i}$ that maximizes the mutual information between the explanation and the prediction $\hat{y}_{i}$. Formally, GNNExplainer optimizes the following objective:
(5) $\max_{\mathcal{G}_{i}^{\text{XAI}},\,X_{i}^{\text{XAI}}}\;\mathbb{MI}\left(\hat{y}_{i};\mathcal{G}_{i}^{\text{XAI}},X_{i}^{\text{XAI}}\right),$
@@
4. Results
4.1. Evaluation Metrics
For model training, we utilized the Adam optimizer with a learning rate of 0.01 and a weight decay of $1e^{-6}$. We also set the number of epochs to 100. We employed 5-fold nested cross validation, in which the train set is split into train/validation for grid search optimization. Details of the grid search parameter space are shown in Supplementary Table 3 and Supplementary Table 4.
For model evaluations, we employed area under the receiver operator curve (AUC) and F1 score (Rainio et al., 2024). For benchmarking and in our ablation studies, both metrics were calculated across all phenotypes. For evaluating our Hierarchical Network Classification model on specific atherosclerosis phenotypes, only AUC was calculated. In all evaluations, both metrics were calculated across 5-fold stratified cross validation.
4.2. Classification and Profiling Performance